how many milliliters of a 10.0% saline solution are needed to make 500 mL of 0.9% saline solution?
1
Expert's answer
2020-09-24T11:21:12-0400

Let V ml of a  10.0% saline solution  is consider .


So , v* 10 / 100 is the mass of solute .------------------- 1 .


To Make Solution of  500 mL of 0.9% saline solution


So , 500 * 0.9 / 100 is the mass of solute -------------------- 2 .


eqn 1 = 2 .


V* 10 / 100 = 500 * 0.9 / 100


V = ( 500 * 0.9 / 10 ) ml = 45 ml . Answer .